The exponential model A=993.9e^0.006 t describes the population A, of a country in millions, t years after 2003. Use the model to determine when the population of the country will be 1037 million
Answer by stanbon(75887)

The exponential model A=993.9e^0.006 t describes the population A, of a country in millions, t years after 2003. Use the model to determine when the population of the country will be 1037 million
------------------------------------------
993.9*e^(0.006t) = 1037x10^6
---
e^(0.006t) = 1043364.524
---
Take the natural log of both sides:
0.006t = 13.8580
t = 2308.88
------
Year: 2003 + 2309 = 4312
====================================
Cheers,
Stan H.